OpenStudy (amonoconnor):

Find the Differential of: "y = ln[sqrt.(1+x^2)]" Any and all help is greatly appreciated!

OpenStudy (amonoconnor):

I got: \[y= \ln \sqrt{1+x^2}\] \[dy = \frac{1}{(x^2 +1)^{1/2}}*(2x)\] \[= \frac{2x}{(x^2+1)^{1/2}}\] BUT... my book has: \[\frac{x}{(x^2+1)}\]

OpenStudy (amonoconnor):

Can anyone explain where I went wrong?

OpenStudy (amonoconnor):

Specifically, I'm wondering where the power of 1/2 goes in the denominator. I see how I messed up the coefficient of 2 in the numerator, but where do the square roots go?

OpenStudy (amonoconnor):

@freckles